CompanyAandB

StatementofFinancialPosition

AsofDecember31, 2020

-Cash(Company A) - 35,857.50

-Cash(Company B) - 24,387.00

-AccountsRecievable(Comapny A) - 175,000.00

-AccountsRecievable(Company B) - 37,837.50

-Inventory(Company A) - 137,715.00

-Inventory(Comapny B) - 21,945.10

-PrepaidExpenses(Company A) - 10,000.00

-PrepaidExpenses(Company B - 7,500.00

-Equipment,net(Company A) - 573,441.00

-Equipment,net(Comapny B) - 153,700.80

-TotalAssets(Company A) - 932,013.50

-TotalAssets(Comapny B) - 245 370.40

-CurrentLiabilities(Company A) - 71,857.50

-CurrentLiabilities(Company B) - 38,882.80

-LongTermDebts(Company A) - 119,350.00

-LongTermDebts(Company B) - 32,582.00

-Owner'sCapital(Company A) - 740,806.00

-Owner'sCapital(Company B) - 173,905.60

-Totalof(Comapny A) - 932,013.50

-Totalof(Comapny B) - 245,370.40

Answerthefollowing:

1. Compute for theprofitability ratios of CompanyAand B.

2. Computefor operational ratios of companyA and B.

3. Compute for thefinancial health ratios of CompanyA and B.


1
Expert's answer
2021-10-20T10:06:03-0400

Solution:

1.). Profitability ratios:

Current ratio = "\frac{Current\;Assets}{Current\;Liabilities}"


Company A = "\frac{358,572.5}{71,857.50} = 5"


Company B = "\frac{91,669.6}{38,882.80} = 2.4"

 

2.). Operational ratios:

Debt to Asset ratio = "\frac{Total\;Debt}{Total\;Assets}"


Company A = "\frac{119,350}{932,013.50} = 0.13"


Company B = "\frac{32,582}{245,370.40} = 0.13"

 

Debt to Equity ratio = "\frac{Total\;Debt}{Shareholders\;Equity }"


Company A = "\frac{119,350}{740,806} = 0.16"


Company B = "\frac{32,582}{173,905.60} = 0.19"

 

3.). Financial health ratios:

Quick ratio = "\frac{Cash +Cash\;Equivalents + Accounts\;Receivable}{Current\;Liabilities }"


Company A "=\frac{210,857.5}{71,857.5} = 2.93"


Company B = "\frac{62,224.5}{38,882.80} = 1.60"

 

Cash ratio = "\frac{Cash +Cash\;Equivalents }{Current\;Liabilities }"


Company A = "=\frac{35,857.5}{71,857.5} = 0.5"


Company B = "\frac{24,387}{38,882.80} = 1.63"